### v4

# Digital Simulation of Non-Abelian Parafermions in Superconducting Circuits

Abstract
Parafermions, which can be viewed as a fractionalized version of Majorana modes, exhibit profound non-Abelian statistics and emerge in topologically ordered systems, while their realization in experiment has been challenging. Here we propose an experimental scheme for the digital simulation of parafermions and their non-Abelian braiding statistics in superconducting (SC) circuits by realizing the $\mathbb{Z}_{d}$ plaquette model on a two-dimensional lattice. Two protocols using quantum circuits and non-destructive measurements are introduced to prepare the ground state, on which parafermion pairs are created by engineering dislocations. We then propose a generalized code deformation approach to realize the fusion and non-Abelian braiding of parafermion modes, and show the application of this approach to the $\mathbb{Z}_{3}$ parafermions. We also examine the experimental parameter regime to confirm the feasibility of our scheme in SC devices. This work extends previous quantum simulation of twist defects in SC qubits to qudit systems, and opens up a way for parafermion-based high-dimensional topological quantum computing with experimental feasibility.
I Introduction
The recent rapid advancement in quantum processors enables simulation of many-body quantum states and exploration of the intriguing physics they exhibit [1, 2, 3, 4, 5, 6]. This provides a way to study and manipulate exotic particles that have yet to be realized or are difficult to manipulate in real materials [7, 8, 9, 10, 11]. A notable example is the simulation of anyons and topological defects in topologically ordered states (topological orders) [3, 4, 12, 13, 14, 15, 16, 17]. In particular, recent progress in digital simulation of twist defects in the $\mathbb{Z}_{2}$ toric code model with superconducting (SC) qubits demonstrates the (projective) non-Abelian braiding statistics [3, 4]. These efforts advance topologically protected quantum information processing, providing a foundation for the potential realization of topological quantum computing.
Twist defects in the $\mathbb{Z}_{2}$ toric code model exhibit non-Abelian statistics of Majorana fermion modes, which is described by the Ising anyon model [18, 19]. Quantum computing schemes based on Majorana modes have been widely studied, but their experimental realization in one- or two-dimensional condensed matter systems faces significant challenges [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. More critically, braiding of Majorana modes alone is insufficient for universal quantum computing [49, 50]. Universal quantum computing schemes based on Majorana modes require additional non-topological assistance methods, making them not strictly topological [51, 52]. $\mathbb{Z}_{d}$ parafermions generalize Majorana modes (which are a special case for $d=2$) and exhibit more complex and nonlinear non-Abelian braiding statistics [53, 54]. In contrast to Majorana modes, parafermions can encode multi-level computational units, or qudits, which offer a larger state space for storing and processing information, as well as advantages in reducing circuit complexity and improving algorithm efficiency [55]. Braiding of parafermions allows the generation of the single-qudit Clifford group for any $d$, and the many-qudit Clifford group for odd $d$ [56]. Fault-tolerant non-Clifford gates can be performed by magic state distillation with high fidelity [57, 58, 59].
Proposals for realizing parafermions typically involve systems with strong electron-electron interactions, which however pose serious experimental challenges [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. On the other hand, to engineer the topological defects is an alternative general approach for realizing non-Abelian statistics within Abelian topological orders [76]. This provides a simpler way to realize parafermion defects in Abelian topologically ordered states. Furthermore, the braiding of Abelian anyons of the underlying state facilitates performing topological non-Clifford gates. A typical class of stabilizer Hamiltonian models that realize Abelian topological orders is the $\mathbb{Z}_{d}$ plaquette models with qudits on a two-dimensional square lattice, in which certain lattice dislocations realize parafermions [77]. As a stabilizer model, the digital simulation of its ground state and excited states with a few quasiparticles is in principle achievable on experimental platforms with coherent control of qudits. The SC devices, such as transmons, which possess readily addressable higher energy levels, constitute promising candidates for implementing such platforms [78]. Recent works on SC devices have demonstrated high-fidelity single- and two-qutrit (qudit with $d=3$) entangling operations [79, 80, 81, 82, 83, 84, 85]. Nevertheless, realization of the parafermion modes and implementation of their non-Abelian statistics necessitate qualitatively more complicated quantum gate operators and involve addressing non-trivial physical challenges in comparison with the Majorana modes [4, 3], and it remains unclear how to achieve such quantum simulation in the real SC circuits. More precisely, while the generalization from Majorana mdoes to parafermion modes may appear straightforward at the level of mathematical model construction, its physical realization remains significantly non-trivial. First, the fundamental single- and two-qudit gates are unitary but not Hermitian, requiring careful consideration of both the operations and their Hermitian conjugates in circuit implementations. Second, the multi-level structure of qudits necessitates novel physical processes to control level transitions and to implement entangling operations involving higher levels. Examples include the two-photon process for controlling the 00-$2$ transition in qutrits [82] to realize high-fidelity generalized Hadamard gate, and implementing control gates based on tunable cross-Kerr interactions [81, 84]. Third, parafermions exhibit richer non-Abelian braiding statistics and fusion rules compared to Majorana modes, making it physically non-trivial to identify suitable observables for characterizing their braiding and fusion properties. This work aims to address these challenges and integrate recently developed qudit techniques in SC devices to develop an experimentally feasible scheme for simulating parafermions and their fusion and braiding, providing insights into the parafermion-based high-dimensional topological quantum computing.
Here, we summarize the main results of our work. In this work, we propose an experimentally feasible scheme to simulate parafermions and their fusion and non-Abelian braiding in SC circuits, focusing on the $\mathbb{Z}_{d}$ plaquette model on a planar surface. We introduce two efficient ground state preparation protocols: one using quantum circuits and the other employing non-destructive measurements. By engineering dislocations in the lattice, pairs of parafermions are created. With a novel generalized code deformation approach, as proposed in the present work, these parafermions are moved on the planar surface, enabling a feasible way for their fusion and braiding operations. The concrete example of $d=3$ parafermions is studied in detail. The real parameter conditions are examined, with the feasibility of our scheme in experiment being confirmed. We note that while the present simulation is designed for the SC devices, it is also compatible with other experimental platforms capable of implementing the gates that we use. Our work provides a theoretical foundation for the imminent digital simulation of parafermions in quantum computing platforms, and shall advance topologically protected high-dimensional quantum information processing.
The remainder of the paper is organized as follows. In Sec. II, we introduce the $\mathbb{Z}_{d}$ plaquette model, its elementary excitations, and parafermion defects. In Sec. III, we present the protocol for ground state preparation and parafermion creation. Sections IV and V outline the schemes for realizing the fusion and non-Abelian braiding of parafermions. In Sec. VI, we discuss the experimental parameters of the SC devices for realizing our scheme. Finally, in Sec. VII, we summarize our findings and discuss potential applications in high-dimensional topological quantum computing.
II The Model
Figure 1: Illustration of the model with a pair of prafermions. (a) The $\mathbb{Z}_{d}$ plaquette model on a checkerboard lattice with a dislocation colored by light blue, where the small white dots represent qudits. The red cross marks the defect associated with the dislocation, identified as the parafermion mode $\gamma$. The red dotted line indicates the branch cut connecting the two parafermions. The arrowed dark blue curves represent string operators that create $e-\overline{e}$ and $m-\overline{m}$ excitations, which reside on the light and dark plaquettes, respectively. In our convention, the $e$ or $m$ particle is located at the start of the string operator and its dual is located at the end of the string operator. Stabilizers $O_{p}$ on both dark and light plaquettes are shown. Graphical representations of the generalized Pauli operators, commutation relation, and stabilizers are shown. (b) An illustration of the energy diagram and microwave driving of a typical transmon as a multilevel system. The frequency difference $\omega_{12}-\omega_{01}$ characterizes the anharmonicity. (c) Dislocations are engineered with two neighboring stabilizers being replaced with a single one. Applying a two-qudit quantum circuit moves and separates the parafermion defects. (d) Schematic of the cross-Kerr interaction that is used to realize two-qutrit gates. (e) A configuration that realizes the tunable cross-Kerr interaction using a frequency-tunable coupler, as demonstrated in Ref. [81].
We start by considering the $\mathbb{Z}_{d}$ plaquette model on the two-dimensional square lattice with a checkerboard pattern [Fig. 1(a)]. On each vertex $i$, there is a qudit with $d$ basis states $\left|{m_{i}}\right\rangle$ ($m_{i}=0,1,2...,d-1$). Physically, the qudit can be realized by the intrinsic multi-dimensional Hilbert space of SC devices, such as transmons. Their coherence properties and sufficient anharmonicity enable to function as qudits, with the lowest $d$ energy levels serving as the qudit basis states [Fig. 1(b)]. The Hilbert space is spanned by the qudit states on all the vertices. The Hamiltonian is:
$H=-\sum_{p}O_{p}+h.c.,\,O_{p}=Z_{1}X_{2}Z_{3}^{\dagger}X_{4}^{\dagger},$ (1)
where the labels of the qudits are shown in the top of Fig. 1(a), and the sum of plaquettes $p$ is for all the light and dark plaquettes. Here, $Z$, $X$, $Z^{\dagger}$, and $X^{\dagger}$ are generalized Pauli operators for qudits. They are defined by the properties: $X^{d}=\mathbbm{1},\,Z^{d}=\mathbbm{1},\,XZ=\omega ZX,$ where $\omega=e^{\frac{2\pi i}{d}}$ and act on the qudit basis as $Z\left|{m}\right\rangle=\omega^{m}\left|{m}\right\rangle$ and $X\left|{m}\right\rangle=\left|{m-1\,(\text{mod}\quad d)}\right\rangle$. The experimental realization of single-qudit operations can be achieved through microwave driving resonant with specific qudit state subspaces. For $d=3$ (i.e., a qutrit), any single-qutrit gate can be decomposed into a sequence of operations acting on two-level subspaces [86, 83], and implemented with high fidelity by accounting for the AC Stark effect arising from anharmonicity [86, 87]. It is convenient to introduce a graphical representation [Fig. 1(a)] for the generalized Pauli operators and their commutation relations. With this representation, the stabilizer $O_{p}$ is described by a loop operator around the plaquette $p$, which satisfies $O_{p}^{d}=1$ and has eigenvalues $1$, $\omega$,…, $\omega^{d-1}$. Stabilizers on different plaquettes commute with each other; therefore, the Hamiltonian Eq. (1) is exactly solvable. The ground state $\left|{\Psi_{0}}\right\rangle$ satisfies:
$O_{p}\left|{\Psi_{0}}\right\rangle=O_{p}^{\dagger}\left|{\Psi_{0}}\right\rangle=\left|{\Psi_{0}}\right\rangle,\,\forall p,$ (2)
and the ground-state projector is:
$\Pi^{0}=\prod_{p}\Pi_{p}^{0},\quad\text{where}\quad\Pi_{p}^{0}=\frac{1}{3}\left(\mathbbm{1}+O_{p}+O^{\dagger}_{p}\right).$ (3)
The model defined by Eq. (1) provides a Hamiltonian realization of the topological gauge theory with gauge group $\mathbb{Z}_{d}$. The stabilizers on light (dark) plaquettes correspond to local Gauss’s law (flatness condition) constraints. The ground state, defined by Eq. (2), is the state where all local constraints are satisfied, while excited states arise from violations of these constraints. For $d=2$, the stabilizer has only two eigenvalues $1$ and $-1$, indicating whether a local constraint is satisfied or violated. The elementary excitations—violations of Gauss’s law and flatness condition—are termed charge ($e$) and flux ($m$) excitations, respectively, following its physical meaning in gauge theory. In the $\mathbb{Z}_{2}$ gauge theory, these excitations are self-dual, satisfying $e^{2}=1$ and $m^{2}=1$. For $d\geq 3$, however, charge and flux excitations are no longer self-dual and can appear in multiple powers. Specifically, the elementary excitations of Eq. (1) include charge excitations $e,e^{2},...,e^{d-1}$ on light plaquettes and flux excitations $m,m^{2},...,m^{d-1}$ on dark plaquettes [see Appendix. A for an algebraic characterization of these excitations]. More general excitations involve composites $e^{k}m^{d-k}$, known as dyons. When a charge excitation $e$ encircles a flux excitation $m$ counterclockwise, or vice versa, the system acquires a phase factor $\overline{\omega}\equiv\omega^{d-1}$ due to the Aharonov-Bohm effect. Consequently, charge, flux, and dyon excitations are Abelian anyons, created by string operators [graphically represented in Fig. 1(a)]. The ground state of Eq. (1) exhibits the Abelian topological order characterized by the quantum double $D(\mathbb{Z}_{d})$ [88].
To incorporate non-Abelian statistics with the Abelian topological order model, we need to engineer dislocations. The dislocation in the square lattice that we consider is an extended object formed by tilting the edges along a line [the light blue region in Fig. 1(a)]. Associated with the dislocation are two point-like defects characterized by trivalent vertices on the pentagon plaquettes. In the $\mathbb{Z}_{d}$ plaquette model, the dislocation is associated with deforming stabilizers $O_{p}$ and $O_{p^{\prime}}$ on two neighbouring plaquettes to a single stabilizer $O_{p^{\prime}}O_{p}$ [Fig. 1(c)]. This deformation effectively removes the edge between the plaquettes and reduces the number of stabilizers by one. Therefore, $N$ such dislocations result in the $Nd$-fold degeneracy, with each defect having a quantum dimension $\sqrt{d}$ that indicates the non-Abelian nature of the defect. These non-Abelian defects manifest parafermionic behavior, whose underlying physics is the defect-induced electric-magnetic exchange symmetry. To illustrate this, consider creating an $e-\overline{e}$ charge pair by applying a $Z$ gate to a qudit, and moving the $\overline{e}$ particle through the branch cut connecting two defects using a string operator [Fig. 1(a)]. The $\overline{e}$ excitation transforms into the $\overline{m}$ excitation, indicating an $e-m$ exchange. This shows that the defect identifies the dyon exictation $e^{k}m^{d-k}$ ($k=1,2,...,d-1$) with the trivial topological charge $1$, giving rise to the following fusion rules:
$\gamma\times\gamma=1+\sum_{k=1}^{d-1}e^{k}m^{d-k},$ (4)
where $\gamma$ denotes the defects. Equation (4) is exactly the fusion rules of parafermions. For clarity, we henceforth refer to these defects as parafermions. By applying certain two-qudit operations, whose circuit realizations are shown later, we can move the parafermions apart in space. This allows to realize the fusion and braiding of parafermions. For convenience, we focus on the qutrit case, i.e., the $\mathbb{Z}_{3}$ plaquette model in the following. Generalizing the methods to larger $d$ is straightforward.
III Ground state preparation and creation of parafermions
Figure 2: Illustration of the ground state preparation and parafermion creation protocol. (a) The $\mathbb{Z}_{3}$ plaquette model with a $6\times 6$ array of qutrits. (b) Explicit procudure for the ground state preparation. The system is initialized with all qutrits in $\left|{0}\right\rangle$. Red dots indicate the representative qutrits used to prepare the light plaquettes in the ground state, while green dots represent the control qutrits used to apply controlled two-qutrit gates for preparing the dark plaquettes in the ground state. The plaquettes in the ground state are colored pale yellow for light plaquettes and gold for dark plaquettes. (c) Applying the generalized Hadamard gate to the representative qutrits prepares the light plaquettes in the ground state. (d) A circuit is applied to the dark plaquettes to prepare them in the ground state. (e) Circuits for preparing boundary dark plauettes in the ground state. (f) Measuring two neighboring qutrits (encircled by dashed line) creates a pair of parafermions. (g) An alternative method for creating parafermions involves ground state preparation on a lattice with a hole, represented by a light blue region.
In this section, we present methods to efficiently prepare the ground state of the stabilizer Hamiltonian in Eq. (1) on a planar surface. With the ground state, we show how parafermions can be created. For illustration, we focus on the case of $d=3$, where the fundamental degrees of freedom are qutrits.
III.1 Ground state preparation
We consider the $\mathbb{Z}_{3}$ plaquette model with $6\times 6$ qutrits [Fig. 2(a)]. The system is initially in the product state $\left|{0}\right\rangle^{\otimes 36}$. The ground state of the model is long-range entangled, which requires two-qutrit entangled operations for the preparation. This can be experimentally achieved through the tunable cross-Kerr interactions between two qutrits [Fig. 1(d)]. A specific configuration for this tunable coupling is shown in Fig. 1(e), where nearest-neighbor transmons are coupled via frequency-tunable couplers. Adjusting the frequency of the coupler modifies the strength of the cross-Kerr interaction, which, in combination with single-qutrit operations, enables the realization of efficient conditional phase gates [81]. Alternatively, the conditional phase gate can also be dynamically realized by the differential AC Stark shift on two transmon qutrits with static coupling [84].
To prepare the ground state $\left|{\Psi_{0}}\right\rangle$ in Eq. (2), we first select a set of representative qutrits on light plaquettes [red dots in Fig. 2(b)] and apply the generalized Hadamard gate to them. The generalized Hadamard gate is defined by the quantum Fourier transform [55, 89]:
$H^{(d)}\left|{j}\right\rangle=\frac{1}{\sqrt{3}}\sum_{i=1}^{d-1}\omega^{ij}\left|{i}\right\rangle.$ (5)
In the case of qutrits, the generalized Hadamard gate acts on the state $\left|{0}\right\rangle$ as:
$\left|{+}\right\rangle:=H^{(3)}\left|{0}\right\rangle=\frac{1}{\sqrt{3}}\left(\left|{0}\right\rangle+\left|{1}\right\rangle+\left|{2}\right\rangle\right),$ (6)
which is the common eigenstate of $X$ and $X^{\dagger}$ with eigenvalue $1$. Experimentally, the generalized Hadamard gate can be implemented by decomposing the gate into time evolutions as $H^{(3)}=e^{-iH_{d}t}e^{-iH_{o}t}$, where $H_{o}=\sum_{i<j}m_{ij}\left|{i}\right\rangle\langle{j}|+h.c.$ is off-diagonal, and $H_{d}=\text{diag}(\phi_{0},\phi_{1},\phi_{2})$ is diagonal with complex parameters $m_{ij}$ and real parameters $\phi_{i}$. The generator $H_{o}$ is experimentally realized by simultaneously microwave driving transitions between the three pairs of energy levels of the qutrit, while $H_{d}$ is implemented by shifting the phases of drive fields in the resonant control pulse [90].
After applying the $H^{(3)}$ to each representative qutrit, all the light plaquettes (including the light boundary plaquettes) are prepared in the ground state [Fig. 2(c)]. Next, we select the control qutrit on each dark bulk plaquette [green dots in Fig. 2(b)] and apply $CZ$, $CZ^{\dagger}$, and $CX$ gates to prepare all the dark plaquettes in the ground state [Figs. 2(d) and 2(e)]. The controlled two-qutrit gates must be applied layer by layer, ensuring that the state of the representative qutrit remains unchanged until the controlled operations on their respective plaquettes are applied. Ground state preparation on the dark plaquettes does not cause any light plaquette to deviate from the ground state, as their stabilizers commute with each other. In this way, we prepare the entire system in the ground state. The depth of the ground state preparation circuit depends linearly on the size of the lattice ($\sqrt{n}$ in our set up, with $n$ the number of qutrits).
[…]
$\left|{\Psi_{tot}}\right\rangle=\frac{1}{\sqrt{3}}\left(\left|{\tilde{0}}\right\rangle+\left|{\tilde{1}}\right\rangle+\left|{\tilde{2}}\right\rangle\right)\otimes\left|{\Psi_{p}}\right\rangle,$ (7)
where the $\left|{\tilde{j}}\right\rangle$ denotes the ancillary state, and $\left|{\Psi_{p}}\right\rangle$ denotes the plaquette state. Then we act $O_{p}$ conditioned on the ancillary state, transforming the state as
$\tilde{C}O_{p}\left|{\Psi_{tot}}\right\rangle=\frac{1}{\sqrt{3}}\left(\left|{\tilde{0}}\right\rangle\left|{\Psi_{p}}\right\rangle+\left|{\tilde{1}}\right\rangle O_{p}\left|{\Psi_{p}}\right\rangle+\left|{\tilde{2}}\right\rangle O_{p}^{\dagger}\left|{\Psi_{p}}\right\rangle\right),$ (8)
where tilde indicates that the conditional gate $\tilde{C}O_{P}$ is controlled by the ancillary qutrit. The gate $\tilde{C}O_{P}$ consists of $\tilde{C}Z_{1}$, $\tilde{C}X_{2}$, $\tilde{C}Z_{3}^{\dagger}$, and $\tilde{C}X_{4}^{\dagger}$ gates, where the subscripts denote the order of qutrits in a stabilizer, as defined in Eq. (1). After applying a $H^{(3)}$ gate to each ancillary qutrit, we measure all the ancillary states at the same time to project each of them onto $\left|{\tilde{0}}\right\rangle$. The plaquette state transforms as
$\tilde{P}(0)\tilde{H}^{(3)}\tilde{C}O_{p}\left|{\Psi_{tot}}\right\rangle$ $=$ $\frac{1}{3}\left(\mathbbm{1}+O_{p}+O^{\dagger}_{p}\right)\left|{\Psi_{p}}\right\rangle$ (9)
$=$ $\Pi_{p}^{0}\left|{\Psi_{p}}\right\rangle,$
where $\tilde{H}^{(3)}$ indicates that the operator acts on the ancillary qutrit, and $\tilde{P}(0)$ is a projector that projects the ancillary qutrit onto the $\left|{\tilde{0}}\right\rangle$ state. Therefore, after the procedure described in Eqs. (7), (8), and (9), each plaquette is subjected to a ground-state projector and prepared in the ground state. This method can be directly generalized to the more general qudit cases.
Fidelity of the ground state preparation can be obtained by the standard quantum state tomography (QST) method to read out the density matrix [91]. With the density matrix, the expectation value $\langle\Pi_{p}^{0}\rangle$ of the ground state projector on each plaquette can be computed. By repeating the experiment multiple times, one obtains the average value $\overline{\langle\Pi_{p}^{0}\rangle}$ for each plaquette. If $\overline{\langle\Pi_{p}^{0}\rangle}=1$, it indicates perfect preparation of the ground state on the plaquette $p$; if $\overline{\langle\Pi_{p}^{0}\rangle}<1$, it indicates that the plaquette deviates from the ground state. This is the typical method used to characterize ground state preparation in digital simulations with SC qubits, and it is extended to the qutrit case in this work.
III.2 Creation of parafermions
With the ground state prepared, we now proceed to create a pair of parafermions. To achieve this, we select two neighboring qutrits and measure them in the $XZ$ basis if they are vertically adjacent or in the $XZ^{\dagger}$ basis if they are horizontally adjacent. By measurement, we project the qutrits onto the eigenstate corresponding to the eigenvalue $1$ of $XZ$ or $XZ^{\dagger}$: $\frac{1}{\sqrt{3}}(\left|{0}\right\rangle+\overline{\omega}\left|{1}\right\rangle+\left|{2}\right\rangle)$ for $XZ$, and $\frac{1}{\sqrt{3}}(\left|{0}\right\rangle+\omega\left|{1}\right\rangle+\left|{2}\right\rangle)$ for $XZ^{\dagger}$. Since $XZ$ and $XZ^{\dagger}$ are creation operators of $e\overline{m}-\overline{e}m$ excitation pairs, this projection projects the qutrits in the coherent state of these creation operators, effectively condensing the $e\overline{m}$ and $\overline{e}m$ excitations in the local region [the light blue region in Fig. 2(f)]. Because the two qutrits are projected separately, they are not directly entangled, which effectively removes the edge between them, thereby creating a pair of parafermions in the condensate.
An alternative method arises in the presence of natural holes in the lattice [Fig. 2(g)], which result from imperfections in the geometric arrangement of qutrits due to fabrication constraints in chip manufacturing. Exploiting this imperfection, a pair of parafermions can be created by preparing the ground state on the imperfect lattice. However, for the region containing the hole, additional controlled gates, such as the controlled-$XZ$ gate, are required to enforce the ground-state condition. An example featuring a minimal hole is illustrated in Fig. 2(g).
IV FUSION OPERATIONs OF THE GENERATED PARAFERMIONS
In this section, we propose the scheme to realize the fusion of parafermions. The previous section presents a method for creating parafermion pairs through engineering dislocations. Nevertheless, since the parafermions remain adjacent, studying the fusion rules of an individual parafermion is difficult because the nearby parafermion can influence the fusion and introduce ambiguity in the interpretation of results. To address this, we generalize the code deformation method to the qudit setting, deriving the operators that facilitate the spatial movement of parafermions. This generalization is technically non-trivial due to the non-Hermitian nature of the stabilizers, requiring careful treatment of both the stabilizer and its Hermitian conjugate. Physically, this property is intrinsically linked to the non-self-duality of anyons and the nonlinearity of parafermion operator transformations, a connection we leave for future work. The generalized code deformation method yields operators that enable the horizontal and vertical movement of parafermions along the lattice. This capability serves as a powerful tool for parafermion manipulation and further facilitates essential operations, such as braiding [see Sec. V]. Given its significance and fundamental role in this study, we first introduce the generalized code deformation method before combining it with string operators to establish a protocol for realizing the fusion of parafermions.
IV.1 Generalized code deformation
Code deformation is a method used to convert one error-correcting code into another by modifying the stabilizers, enabling fault-tolerant logical gate operations within surface codes [92, 93]. Beyond quantum information science, code deformation also finds applications in condensed matter physics. Topological surface codes, such as the toric code, are rooted in topological quantum field theory, which provides the framework for topological orders. In this context, code deformation involves altering the topological structure of the field theory, such as engineering topological defects, to enrich the properties of topological orders [94]. While the method has been extensively studied in qubit systems, its application to qudit systems remains relatively unexplored. In this subsection, we generalize the code deformation method to the qudit system. This generalization enables the derivation of unitary operators that facilitate the spatial movement of parafermions across the lattice. Our work extends the applicability of code deformation to a broader class of qudit surface codes.
Given a qudit stabilizer $O_{p}$ on the plaquette $p$ and Hamiltonian $H=-\sum_{p}O_{p}+h.c.$, the local ground state $\left|{\psi}\right\rangle$ satisfies $O_{p}\left|{\psi}\right\rangle=O_{p}^{\dagger}\left|{\psi}\right\rangle=\left|{\psi}\right\rangle$. After a code deformation, such as a change in the geometry of the lattice, the stabilize changes to $O^{\prime}_{p}$, and the corresponding local ground state is $\left|{\psi^{\prime}}\right\rangle$. The new local ground state is related to the original one by a unitary operator $U$: $\left|{\psi^{\prime}}\right\rangle=U\left|{\psi}\right\rangle$. For qutrits, the unitary operator is given by
$U=\frac{1}{\sqrt{3}}\left(\mathbbm{1}+O^{\prime\dagger}_{p}O_{p}+O^{\prime}_{p}O_{p}^{\dagger}\right).$ (10)
For our purposes, we study the code deformations depicted in Fig. 3(a) and Fig. 3(b), where the former moves the parafermion vertically and the latter moves the parafermon horizontally. The associated unitary operator for the vertical move is:
$U_{v}(i,j)=\frac{1}{\sqrt{3}}\left(\mathbbm{1}+\omega Z_{i}^{\dagger}X_{j}+Z_{i}X_{j}^{\dagger}\right),$ (11)
and for the horzontal move:
$U_{h}(i,j)=\frac{1}{\sqrt{3}}\left(\mathbbm{1}+\omega Z_{j}X_{i}+Z_{j}^{\dagger}X_{i}^{\dagger}\right),$ (12)
where $i$ and $j$ denote the initial and final positions of the parafermion, respectively. Both $U_{v}$ and $U_{h}$ are two-qutrit operators, and their matrices in the basis $\{\left|{00}\right\rangle,\left|{01}\right\rangle,\left|{02}\right\rangle,\left|{10}\right\rangle,\left|{11}\right\rangle,\left|{12}\right\rangle,\left|{20}\right\rangle,\left|{21}\right\rangle,\left|{22}\right\rangle\}$ are:
Figure 3: Generalized code deformation and the fusion of parafermions. (a) Code deformation that moves the parafermion vertically. (b) Code deformation that moves the parafermion horizontally. (c) Illustration of the creation of a composite excitation and braiding $e$ or $\overline{e}$ with one parafermion. After the braiding, the excitations fuse to the trivial topological charge. (d) Explicit procedure for realizing the fusion of parafermions. Observables are characterized by $\langle\Pi_{p^{\prime}}^{d-1}\rangle=\langle\Pi_{p}^{1}\rangle=1$. (e) Quantum circuit and two-qutrit gates used in the procedure.
$U_{v}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&\omega&1&0&0&0&0&0&0\\ 1&1&\omega&0&0&0&0&0&0\\ \omega&1&1&0&0&0&0&0&0\\ 0&0&0&1&1&\omega&0&0&0\\ 0&0&0&\omega&1&1&0&0&0\\ 0&0&0&1&\omega&1&0&0&0\\ 0&0&0&0&0&0&1&\overline{\omega}&\overline{\omega}\\ 0&0&0&0&0&0&\overline{\omega}&1&\overline{\omega}\\ 0&0&0&0&0&0&\overline{\omega}&\overline{\omega}&1\\ \end{pmatrix},$ (13)
and
$U_{h}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&0&0&\omega&0&0&1&0&0\\ 0&1&0&0&\overline{\omega}&0&0&\overline{\omega}&0\\ 0&0&1&0&0&1&0&0&\omega\\ 1&0&0&1&0&0&\omega&0&0\\ 0&\overline{\omega}&0&0&1&0&0&\overline{\omega}&0\\ 0&0&\omega&0&0&1&0&0&1\\ \omega&0&0&1&0&0&1&0&0\\ 0&\overline{\omega}&0&0&\overline{\omega}&0&0&1&0\\ 0&0&1&0&0&\omega&0&0&1\\ \end{pmatrix}.$ (14)
Before proceeding to the realization of the operators $U_{h}$ and $U_{v}$ using quantum circuits, we briefly review the code deformation method used to move Majorana modes in stabilizer codes with SC qubits [3, 4] and compare it with our generalization. In that case, both the original stabilizer $S_{\text{old}}$ and the transformed stabilizer $S_{\text{new}}$ after moving the Majorana mode are Hermitian, simplifying the construction of a unitary operator that transforms the system into an eigenvalue-$1$ eigenstate of the new stabilizer. Specifically, since $S_{\text{old}}$ anticommutes with $S_{\text{new}}$, their product $iS_{\text{new}}S_{\text{old}}$ is Hermitian, motivating the ansatz of the unitary operator $e^{\theta S_{\text{new}}S_{\text{old}}}$, where the parameter $\theta=\frac{\pi}{4}$ is fixed by applying the operator to the eigenvalue-$1$ eigenstate of $S_{\text{old}}$ and requiring that the resulting state be the eigenvalue-$1$ eigenstate of $S_{\text{new}}$. Nevertheless, in the qudit case, this ansatz does not apply. The absence of Hermitian stabilizers complicates the construction of the unitary operator in Eq. (10), making the generalization significantly non-trivial.
The two-qutrit operators $U_{h}$ and $U_{v}$ belong to the Clifford group, meaning they can be realized with a universal set of qutrit gates. The specific circuit for realizing $U_{v}$ operator is shown in Fig. 4(a), where $G_{v}$ is a single-qutrit operator decomposable into rotations within the $\{\left|{0}\right\rangle,\left|{1}\right\rangle\}$ and $\{\left|{1}\right\rangle,\left|{2}\right\rangle\}$ subspaces, along with diagonal phase matrices:
$G_{v}=$ $\begin{pmatrix}1&0&0\\ 0&e^{\frac{i\pi}{3}}&0\\ 0&0&e^{\frac{i\pi}{3}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&\text{cos}(\theta_{1})&-\text{sin}(\theta_{1})\\ 0&\text{sin}(\theta_{1})&\text{cos}(\theta_{1})\end{pmatrix}$ (15)
$\begin{pmatrix}\text{cos}(\theta_{2})&-\text{sin}(\theta_{2})&0\\ \text{sin}(\theta_{2})&\text{cos}(\theta_{2})&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&e^{\frac{i4\pi}{3}}&0\\ 0&0&e^{\frac{i5\pi}{6}}\end{pmatrix}$
$\begin{pmatrix}1&0&0\\ 0&\text{cos}(\theta_{1})&-\text{sin}(\theta_{1})\\ 0&\text{sin}(\theta_{1})&\text{cos}(\theta_{1})\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix},$
where $\theta_{1}=\frac{\pi}{4}$ and $\theta_{2}\approx 1.696\pi$. These rotation and diagonal matrices can be experimentally implemented using microwave drives for level transitions and phase shifts of drive pulses [82, 83]. The circuit implementation for $U_{v}$ is not necessarily unique and optimization algorithms can enhance its efficiency [95]. Machine learning methods may offer further potential improvements. Once $U_{v}$ is implemented, the operator $U_{h}$ can be obtained using the circuit in Fig. 4(b), where the qutrit SWAP gate is implemented using the circuit in Fig. 4(c). Here, the $K^{(3)}$ gate is defined by $K^{(3)}\left|{j}\right\rangle=\left|{3-j}\right\rangle$.
Figure 4: (a) Quantum circuit that realizes the $U_{v}$ operator. (b) Quantum circuit that transforms the $U_{v}$ operator into the $U_{h}$ operator. (b) The qutrit SWAP circuit.
IV.2 Realization of fusion operations
Figure 5: Braiding of parafermion modes. (a) The $\mathbb{Z}_{3}$ plaquette model with four parafermions denoted by $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$ labeld by the red crosses. (b) The fusion tree of parafermions. (c) Illustration of the creation of two pairs of parafermions from the ground state and the full braiding between two parafermions. After the braiding, they fuse to two composite excitations denoted by $f$ and $f^{\prime}$, respectively. (d) Explicit procedure for the full braiding of parafermions $\gamma_{2}$ and $\gamma_{3}$. (e) Two-qutrit gates used in the procedure.
We now present the protocol for realizing the fusion of parafermions. The characteristic fusion rules for parafermions are:
$\gamma\times\gamma=1+e\overline{m}+\overline{e}m,$ (16)
which is a special case of Eq. (4). This rule indicates that the parafermion provides the fusion channels that identify the $e\overline{m}$ and $\overline{e}m$ excitations with the trivial topological charge $1$. The related fusion rule is:
$\gamma\times e\overline{m}/\overline{e}m=1,$ (17)
which shows that a parafermion can absorb either the $e\overline{m}$ or $\overline{e}m$ excitations. This can be achieved by braiding an $e$ or $\overline{e}$ excitation with a parafermion [Fig. 3(c)].
In Fig. 3(d), we explicitly illustrate the procedure for realizing the fusions in Eq. (17). The protocol begins by measuring two horizontally adjacent qutrits in the $XZ^{\dagger}$ basis and projecting them onto the eigenvalue-$1$ eigenstate, thereby creating a pair of parafermions. Since the parafermions are created from the ground state, their initial state is in a fixed fusion channel: $\left|{\gamma,\gamma;1}\right\rangle$. We then separate the two parafermions by moving one downward using the operator $U_{v}$. Next, an $e\overline{m}$ dyon excitation is created by applying two Pauli operators. The $e$ excitation is then braided around one parafermion using a string operator, which consists of a sequence of Pauli operators. After braiding, the $e$ excitation is transformed into the $m$ excitation and proceeds to annihilate the $\overline{m}$ excitation. Once this process is completed, we move the parafermion back to a neighboring position and measure the expectation values $\langle\Pi^{d-1}_{p^{\prime}}\rangle$ and $\langle\Pi^{1}_{p}\rangle$ by the QST method [the (vi) step in Fig. 3(d)]. Before fusion, both $\langle\Pi^{d-1}_{p^{\prime}}\rangle$ and $\langle\Pi^{1}_{p}\rangle$ vanish due to the initial fusion channel $\left|{\gamma,\gamma;1}\right\rangle$. After fusion, the parafermon absorbs the $e\overline{m}$ excitation, resulting in the predicted expectation values $\overline{\langle\Pi^{d-1}_{p^{\prime}}\rangle}=1$ and $\overline{\langle\Pi^{1}_{p}\rangle}=1$, averaged over multiple experimental realizations. A similar procedure can be applied to create and braid the $\overline{e}m$ excitation, providing an experimentally accessible observable for realizing the fusion rules in Eq. (17). Notably, reversing steps (iii-iv) in Fig. 3(d) implements the fusion rule in Eq. (16). In Fig. 3(e), we demonstrate the corresponding circuit and two-qutrit operators used in the procedure.
V Braiding of parafermion modes
After realizing the fusion of parafermions, we now turn to the braiding of them. In Sec. IV, we introduce the generalized code deformation method, which provides two unitary operators that move parafermions vertically and horizontally [Figs. 3(a) and 3(b)]. These operations enable the full braiding of parafermions. To characterize the braiding statistics, we define the computational space basis encoded by parafermions. Compared to Majorana modes, the computational space of $\mathbb{Z}_{3}$ parafermions naturally encodes logical qutrits rather than qubits. This distinction arises from the difference in fusion rules between parafermions and Majorana mdoes, with the former having more fusion channels. Within this larger computational space, parafermions exhibit more intricate braiding statistics, making them harder to measure. We identify the expectation value of local ground-state projectors as an observable for characterizing braiding results.
To encode information, a minimum number of four parafermions is required [Fig. 5(a)]. In this configuration, there are only two parity operators, $\Lambda_{1}=\overline{\omega}\gamma_{1}\gamma_{2}^{\dagger}$ and $\Lambda_{3}=\overline{\omega}\gamma_{3}\gamma_{4}^{\dagger}$ with $\gamma$ the parafermion operator [see Appendix. B]. For convenience, the computational space is restricted to the two-dimensional subspace spanned by $\Lambda_{1}\Lambda_{3}=1$. With this restriction, four parafermions are sufficient to encode a logical qutrit. To find the basis of the computational space, we diagonalize the parity operator $\Lambda_{i}$ and obtain the eigenstates $\{\left|{0}\right\rangle_{\Lambda_{i}},\left|{1}\right\rangle_{\Lambda_{i}},\left|{2}\right\rangle_{\Lambda_{i}}\}$. The logical qutrit state is defined by:
$\left|{\overline{j}}\right\rangle=\left|{j}\right\rangle_{\Lambda_{1}}\otimes\left|{3-j}\right\rangle_{\Lambda_{3}}.$ (18)
On the other hand, physically, the basis of the computational space is defined by the fusion tree of the four parafermions [Fig. 5(b)]. The specific transformation between the basis defined by the eigenstates of the parity operators and the fusion basis $\left|{\gamma_{1},\gamma_{2};1/e\overline{m}/\overline{e}m}\right\rangle\otimes\left|{\gamma_{3},\gamma_{4};1/\overline{e}m/e\overline{m}}\right\rangle$ requires selecting two base points on the lattice and constructing the parafermion operators using string operators whose start and end points are at these base points and satisfying the algebra in Eq. (24[18, 3]. Nevertheless, in this work, we do not make this exact transformation by selecting base points. Instead, we take a more direct approach to measure the braiding consequences of the parafermions, which does not rely on the explicit relation between the fusion basis and the logical qutrit basis $\left|{j}\right\rangle_{\Lambda_{1}}\otimes\left|{3-j}\right\rangle_{\Lambda_{3}}$.
In the logical qutrit basis $\{\left|{\overline{j}}\right\rangle\}$, the half-braiding operator $U_{2}$ that exchanges $\gamma_{2}$ and $\gamma_{3}$ is an off-diagonal matrix, exhibiting the non-Abelian statistics. In a specific representation, the matrix elements of $U_{2}$ are given by: $\langle{\overline{i}}|U_{2}\left|{\overline{j}}\right\rangle=\frac{1}{\sqrt{3}}c_{i-j},$ where $c_{j}=\omega^{\frac{j(j+5)}{2}}$ [see Appendix. B]. The full braiding matrix in the logical qutrit basis is:
$U_{2}^{2}=\frac{1}{\sqrt{3}}\begin{pmatrix}i&i&\frac{2+\overline{\omega}}{\sqrt{3}}\\ \frac{2+\overline{\omega}}{\sqrt{3}}&i&i\\ i&\frac{2+\overline{\omega}}{\sqrt{3}}&i\\ \end{pmatrix},$ (19)
and acts on the logical qutrit state $\left|{\overline{0}}\right\rangle$ as:
$U_{2}^{2}\left|{\overline{0}}\right\rangle=\frac{i}{\sqrt{3}}\left|{\overline{0}}\right\rangle+\frac{2+\overline{\omega}}{3}\left|{\overline{1}}\right\rangle+\frac{i}{\sqrt{3}}\left|{\overline{2}}\right\rangle.$ (20)
The key point is that, regardless of how the logical states are identified in terms of the fusion basis, the fusion state corresponding to the trivial charge, $\left|{\gamma_{1},\gamma_{2};1}\right\rangle\otimes\left|{\gamma_{3},\gamma_{4};1}\right\rangle$, can always be identified with the state $\left|{\overline{0}}\right\rangle$. This follows from the physical insight that a state with a total trivial topological charge has parity $1$. Since the parafermions are created from the ground state, their initial state is $\left|{\gamma_{1},\gamma_{2};1}\right\rangle\otimes\left|{\gamma_{3},\gamma_{4};1}\right\rangle\equiv\left|{\overline{0}}\right\rangle$. After a full braiding of $\gamma_{2}$ and $\gamma_{3}$, the state transforms as Eq. (20).
We now move on to the explicit procedure. In Fig. 5(d), we show the implementation of full braiding of parafermions. The process begins by creating four parafermion $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$ from the ground state, arranged in the configuration shown in Fig. 5(a). In the ground state, measuring $\langle\Pi_{p_{i}}^{0}\rangle$ in these plaquettes yields $\langle\Pi_{p_{i}}^{0}\rangle=1$. We then move $\gamma_{3}$ to the bottom, followed by moving $\gamma_{2}$ to the right. After these steps, we move $\gamma_{2}$ back to its original position and finally return $\gamma_{3}$ to its starting point. After braiding, the parafermions $\gamma_{1}$ and $\gamma_{2}$ fuse to form a composite excitation, denoted by $f$, which consists of the components $1$, $\overline{e}m$, and $e\overline{m}$. Similarly, parafermions $\gamma_{3}$ and $\gamma_{4}$ fuse into another composite excitation $f^{\prime}$ [Fig. 5(c)]. In general, $f$ and $f^{\prime}$ are distinct, but they are related by an electric-magnetic transformation, where $e\leftrightarrow m$. According to Eq. (20), the probability of measuring the topologically trivial charge $1$ from the composite excitation $f$ or $f^{\prime}$ is always $\frac{1}{3}$. Therefore, when measuring the expectation value $\langle\Pi_{p_{i}}^{0}\rangle$ using the QST method on the plaquettes that support the dislocation, the averaged result over multiple repeated experiments is given by $\overline{\langle\Pi_{p_{i}}^{0}\rangle}=\frac{1}{3}$. This provides an experimental observable for the braiding of parafermions. In Fig. 5(e), we show the two-qutrit gates used in this procedure.
VI Experimental parameters
In this section, we briefly discuss typical parameters for SC circuits to realize our scheme, particularly for the case of $d=3$. One key parameter in SC circuits is the fidelity of single- and two-qutrit gates. Recent work has demonstrated the realization of the generalized Hadamard gate for qutrits, achieving a high fidelity of $0.992$ [82]. More challenging is the implementation of high-fidelity two-qutrit gates. In recent study Ref. [84], the authors report the realization of $CZ^{\dagger}$ and $CZ$ gates for qutrits, with fidelities of $0.973$ and $0.952$, respectively. Since the fidelity of single-qutrit gates is much higher than that of two-qutrit gates, we will, as a first approximation, neglect errors from single-qutrit gates in the following analysis. Our ground state preparation protocol requires three two-qutrit gates for each dark plaquette. To achieve a ground state preparation fidelity of $0.90$, the fidelity of the two-qutrit gates needs to be above $0.965$, which is very close to the fidelity achievable with existing technology. Achieving higher level of fidelity would require further technological advances in SC devices, highlighting an important avenue for future research in this field. Other typical parameters include the relaxation time $T^{ij}_{1}$ and the dephasing time $T^{ij}_{2}$ between different levels ($ij\in\{01,12,02\}$), which must be sufficiently long to complete the braiding, fusion procedure, and measurement. Recent studies on transmon qutrits report $T^{ij}_{1}$ and $T^{ij}_{2}$ values typically around $10\mu s$ or higher [81, 84]. Considering that the gate times for high precision control are on the order of $10-100ns$ for single-qutrit gates and $100-800ns$ for two-qutrit gates, these parameters suggest that ground state preparation, along with subsequent fusion and braiding procedures, is highly promising and achievable with the near-future technology [81].
In the main text, we consider a $6\times 6$ qutrit system for illustration, though this is not the minimal setup for our scheme. To stduy the fusion rules, a $3\times 4$ qutrit setup is sufficient, while for braiding, a $4\times 4$ setup is adequate. Reducing the number of qutrits in each row (or column) decreases the number of parafermion movement operations by two, saving the operation time by hundreds to thousands of nanoseconds. This reduction significantly enhances the success rate and fidelity of the experiment. The $6\times 6$ example allows for quantum information processing based on more parafermions, such as $8$ parafermions encoding two logical qutrits and their braiding. Additionally, considering braiding on well-designed quasi-one-dimensional configurations can further minimize the required number of qutrits.
VII Discussion and conclusion
Before concluding, we discuss the potential applications in high-dimensional quantum computing. Recent developments highlight several advantages over qudit-based quantum information processing [96] and flexible simulation of quantum dynamics with qudits [97]. Qudit versions of various quantum algorithms have already been developed [98, 99, 89]. High-dimensional quantum computing based on non-Abelian parafermions is inherently fault-tolerant due to topological protection. While realizing parafermions in physical materials remains a great challenge, our digital simulation scheme offers a promising approach to high-dimensional gate-based topological quantum computing in the near future. One exciting direction is to implement qudit quantum algorithms using parafermions braidings. Another challenge is using digital simulation of parafermions to implement the topologically protected $CX$ gate, involving eight parafermions and twelve half-braidings [56].
In summary, we have proposed in this work an experimental scheme for realizing and manipulating non-Abelian parafermions in SC circuits, which should be also applicable to other quantum simulation and quantum computing platforms, like trapped ions and neutral atom arrays which develop fast recently. We have demonstrated efficient protocols for ground state preparation and parafermion creation, and proposed a generalized code deformation approach to achieve the fusion rules and braiding statistics through experimental observables. This work with experimental feasibility paves the way for the digital simulation of parafermions and their braiding statistics in SC circuits.
Acknowledgments
This work was supported by National Key Research and Development Program of China (2021YFA1400900), the National Natural Science Foundation of China (Grants No. 12425401 and No. 12261160368), the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302000), and by Shanghai Municipal Science and Technology Major Project (Grant No. 2019SHZDZX01).
APPENDIX A ELEMENTARY EXCITATIONS OF THE MODEL
In this section, we investigate the elementary excitations of the model Hamiltonian Eq. (1). The $\mathbbm{Z}_{d}$ plaquette model is an exactly solvable Hamiltonian model, of which the spectrum includes a series of elementary excitations. To characterize them, we define the following projectors:
$\Pi_{p}^{\lambda}=\frac{1}{d}\sum_{k=0}^{d-1}\omega^{\lambda(d-k)}O_{p}^{k},$ (21)
where $\lambda=0,1,...,d-1$. These projectors are observables $[\Pi_{p}^{\lambda},H]=0$ and satisfy
[…]
Therefore, their eigenvalues define a set of topological charges $\{\pi_{p}^{0},\pi_{p}^{1},...,\pi_{p}^{\lambda},...,\pi_{p}^{d-1}\}$ for each plaquette $p$. In particular, the ground states have the topological charge $\{1,0,...,0\}$.
Among the excitation states, we identify that states with topological charge $\{0,1,\dots,0\}$ on a given light (dark) plaquette $p$ correspond to a charge (flux) excitation $e$ ($m$) on $p$. Additional quantum numbers $\pi_{p}^{\lambda}=1$ indicate the presence of a charge (flux) excitation $e^{\lambda}$ ($m^{\lambda}$) on the light (dark) plaquette $p$. For convenience, we define $\overline{e}\equiv e^{d-1}$ and $\overline{m}\equiv m^{d-1}$. These elementary excitations can be created in pairs and moved via string operators. More generally, composite excitations are formed from these elementary excitations.
APPENDIX B ALGEBRAIC THEORY OF PARAFERMIONS AND BRAIDING OPERATOR
In this section, we introduce the algebraic theory of $\mathbb{Z}_{3}$ parafermions. For a general $d$, we refer to Ref. [56]. The algebraic relations for $\mathbb{Z}_{3}$ parafermions are given by:
$\gamma_{i}\gamma_{j}=\omega^{\text{sgn}(j-i)}\gamma_{j}\gamma_{i},\quad\gamma_{i}^{3}=\mathbbm{1},$ (24)
for an ordered set $\{i\}$ with $\omega=e^{\frac{2\pi i}{3}}$. To process quantum information, we need to consider $N$ pairs of parafermions and unitary representations of the braid group $B_{2N}$. It turns out that there are six representations of the braid group in terms of the parity operators $\Lambda_{i}$, which are defined as:
$\Lambda_{i}=\overline{\omega}\gamma_{i}\gamma_{i+1}^{\dagger},$ (25)
where $\overline{\omega}$ denotes the complex conjugate of $\omega$. These operators satisfy the following relations:
[…]
The half-braiding operator that exchanges $\gamma_{i}$ and $\gamma_{i+1}$ is given by:
$U_{i}=\frac{1}{\sqrt{3}}\sum_{m=0}^{2}c_{m}\left(\Lambda_{i}\right)^{m}.$ (28)
These operators must satisfy the Yang-Baxter equation:
$U_{i}U_{i+1}U_{i}=U_{i+1}U_{i}U_{i+1}.$ (29)
Solving the equation yields a family of six representations, which differ by the coefficients $c_{m}$:
$c_{m}=\omega^{\pm\frac{m(m+2r+3)}{2}},\,r=0,1,2.$ (30)
In this paper, we focus on a specific representation where $c_{m}=\omega^{\frac{m(m+5)}{2}}$, and the corresponding operator transformations under half-braiding are:
$U_{i}\gamma_{i}U_{i}^{\dagger}$ $=$ $\overline{\omega}\gamma_{i+1},$
$U_{i}\gamma_{i+1}U_{i}^{\dagger}$ $=$ $\gamma_{i}^{\dagger}\left(\gamma_{i+1}\right)^{2}.$ (31)
With this representation, the half-braiding operator $U_{i}$ becomes:
$U_{i}=\frac{1}{\sqrt{3}}\left(1+\overline{\omega}\gamma_{i}\gamma_{i+1}^{\dagger}+\overline{\omega}\gamma_{i+1}\gamma_{i}^{\dagger}\right),$ (32)
and the full braiding operator is $U_{i}^{2}$.
For a pair of $\mathbb{Z}_{3}$ parafermions, the computational basis, which is subject to the constraint $\Lambda_{1}\Lambda_{3}=1$, is given by the eigenstates of the parity operators $\Lambda_{i}$:
$\Lambda_{i}\left|{m}\right\rangle_{\Lambda_{i}}=m\left|{m}\right\rangle_{\Lambda_{i}}.$ (33)
We denote the basis as $\left|{\overline{k}}\right\rangle=\left|{k}\right\rangle_{\Lambda_{1}}\otimes\left|{3-k}\right\rangle_{\Lambda_{3}}$ with $k=0,1,2$. In this basis, the matrix representation of the half-braiding operator $U_{1}$ is diagonal:
$U_{1}=\sum_{k=0}^{2}\tilde{c}_{k}\left|{\overline{k}}\right\rangle\langle{\overline{k}}|,$ (34)
where $\tilde{c}_{k}=\frac{1}{\sqrt{3}}\sum_{j=0}^{2}\omega^{jk}c_{j}$ is the Fourier transform of $c_{m}$. The half-braiding operator $U_{2}$, however, is off-diagonal:
$\langle\overline{k}|U_{2}|\overline{l}\rangle$ $=$ $\langle\overline{k}|F^{\dagger}U_{1}F|\overline{l}\rangle$ (35)
$=$ $\frac{1}{\sqrt{3}}c_{k-l}=\frac{1}{\sqrt{3}}\omega^{\frac{(k-l)(k-l+5)}{2}},$
where $F$ is the Fourier transformation operator defined by $F=\frac{1}{\sqrt{3}}\sum_{k,m=0}^{2}\omega^{km}|\overline{k}\rangle\left\langle\left.\overline{l}\right|\right.$. Finally, we obtain the matrix of $U_{2}^{2}$ in Eq. (19).